OpenStudy (anonymous):

A circle its center at the center of a square with 8 inch sides. Find the area of the square not covered by the circle. Round the answer to the nearest tenth. A) 57.1 in² B) 9.1² C) 2.0² D) 13.7 in²

jimthompson5910 (jim_thompson5910):

is the circle larger than the square? or the other way around?

jimthompson5910 (jim_thompson5910):

oh wait, nvm

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

the square has 8 inch sides, so the area of the square is 8^2 = 8*8 = 64 square inches

jimthompson5910 (jim_thompson5910):

the circle has a radius of 4 since this is half of 8